OpenStudy (anonymous):

Solve for a: 4/a = 100/45

OpenStudy (anonymous):

You have to cross multiply. 4*45 = 100a 180 = 100a a = 1.8

OpenStudy (anonymous):

To check the answer, simply plug in 1.8 as a. Then you have to cross multiply again. 4/1.8 = 100/45 4*45 = 100* 1.8 180 = 180

OpenStudy (anonymous):

I forgot it is 4a/9 =100/45

OpenStudy (anonymous):

Okay. So I told you how to do it, so try and do this one by yourself. I'll be here to help you if you make a mistake.

OpenStudy (anonymous):

The first step is to cross multiply. Multiply 4a by 45 and 9*100.

OpenStudy (anonymous):

900 = 180a

OpenStudy (anonymous):

Good. Now you have to solve for a. To do this, divide 900 by 180. :)

OpenStudy (anonymous):

5

OpenStudy (anonymous):

a = 5, yes! :)

OpenStudy (anonymous):

Okay, here is another. 3/4 = 21/2+d , that goes to 3(2+D) = 4(21)

OpenStudy (anonymous):

Hold on one minute, please.

OpenStudy (anonymous):

You're right with the first step. Now you have to simplify 3(2 + D). To do this, you multiply 3*2 and 3*D.

OpenStudy (anonymous):

6d

OpenStudy (anonymous):

Close, but no. You have to multiply both the number and the variable inside of the quotations. It would be like this: 6 + 3D.

OpenStudy (anonymous):

Would the answer be 2.

OpenStudy (anonymous):

So, the entire equation would look like this: 6 + 3D = 84.

OpenStudy (anonymous):

Whoa, your fast. Hold on.

OpenStudy (anonymous):

Okay. I tried doing it while you said wait a minute.

OpenStudy (anonymous):

Haha, no the answer would not be 2. Sorry. :(

OpenStudy (anonymous):

Alright, let's take this slowly.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

So, we have this: 3(2 + D) = 4(21)

OpenStudy (anonymous):

First multiply 4 and 21.

OpenStudy (anonymous):

84.

OpenStudy (anonymous):

Correct. So now we have: 3(2 + D) = 84.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

6D=84

OpenStudy (anonymous):

The next step is to simplify 3(2 + D). I forgot the term, but you to multiply both the number and the variable inside of the quotations by 3.

OpenStudy (anonymous):

No, it would be. 3*2 + 3*D, which would be: 6 + 3D.

OpenStudy (anonymous):

The mistake you made was that you multiplied 3*2 and added the D.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Do you understand? The correct way to do this is to multiply both the number (2) and the variable (D). This will leave you with 6 + 3D, because 3*2 = 6 and 3*D is 3D.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So now we have 6 + 3D = 84.

OpenStudy (anonymous):

What do you think we have to do next? It's okay if you don't get it correct. :)

OpenStudy (anonymous):

Divide 84 by 6.

OpenStudy (anonymous):

Yes! And the answer is?

OpenStudy (anonymous):

Subtract 84 by six.

OpenStudy (anonymous):

Oh, I'm so sorry. Yes, subtract. Even I'm losing myself here. Haha. And what's the answer?

OpenStudy (anonymous):

78

OpenStudy (anonymous):

Then you divide. And get 26.

OpenStudy (anonymous):

Yes. So 78 = 3D. What's next?

OpenStudy (anonymous):

Yes. Good job!

OpenStudy (anonymous):

Anything else that you need help with? :)

OpenStudy (anonymous):

just go over it with me,.

OpenStudy (anonymous):

3/0.8 = 12/b

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

So, once again, we have to cross multiply.

OpenStudy (anonymous):

Do you think you can do that for me?

OpenStudy (anonymous):

Can I try it by myself, then you tell me if it is right or not?

OpenStudy (anonymous):

Sure thing.

OpenStudy (anonymous):

So far, 3b = 9.6

OpenStudy (anonymous):

3.2

OpenStudy (anonymous):

Perfect!

OpenStudy (anonymous):

Anything else?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Okay, have a nice day.